OpenStudy (anonymous):

reduce to slope-intercept form :5x - 2y =-10

OpenStudy (nincompoop):

solve for y

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

thanks

OpenStudy (nincompoop):

do you not know how to solve for y?

OpenStudy (anonymous):

@SAMGIRL_08 is offline.

OpenStudy (nincompoop):

2x + 10y = 100 solving for y involves isolating y by itself and it can be achieved by 2x - 2x + 10y = 100 - 2x then 10y = 100 - 2x we have 10y but we want just y so we divide both the left hand side (LHS) and the right hand side (RHS) by 10 y = 10 - 1/5 x

OpenStudy (nincompoop):

i hope the example I provided helps

OpenStudy (anonymous):

why 100 ?